Find general solution trol y%+ 12 y²+36 y = te erot
Answer #1

Tweety 124 + 364 = the but T dete - 1 + 12 x + 1 ) = = Auxillacy equaction 3 Libouti Dom - m² +12m +36 = 0 m² & 6m +6m +36=0from eam von P-I I . ücks).€ 9t; eau avess (10 +902 Here 1 96(6112 [D-6) ²+ 12 (0-6) +26) it2 [DP-128+36+120-78+36]*2 2 est 1geriral isolation or بدامه + (وہدری و آنها بهادری + = T = et Cat Caut - logilt)] Aneplease give you thumb up you think this solution is good